A study found that 45% of the embryo transfers performed in assisted reproductive technology (ART) procedures resulted in pregnancies.
Twenty-four percent of the ART pregnancies resulted in multiple births.
(a) Find the probability that a randomly selected embryo transfer resulted in a pregnancy and produced a multiple birth.
(b) Find the probability that a randomly selected embryo transfer that resulted in a pregnancy did not produce a multiple birth.

Answer by ikleyn(52780)

.
A study found that 45% of the embryo transfers performed in assisted reproductive technology (ART) procedures resulted in pregnancies.
Twenty-four percent of the ART pregnancies resulted in multiple births.


(a)   Find the probability that a randomly selected embryo transfer resulted in a pregnancy and produced a multiple birth.

        P(A & B) = P(A)*P(B) = 0.45*0.24 = 0.108 = 10.8%.


(b)   Find the probability that a randomly selected embryo transfer that resulted in a pregnancy did not produce a multiple birth.

        This probability is  1 - P(B) = 1 - 0.24 = 0.76 = 76%.